1.A sample consists of the following n = 5 scores: 8, 4, 10, 0, 3. (2 pts)

Compute the mean and standard deviation for the sample.

b)Find the z-score for each score in the population from exampl
Answer #1
Values ( X ) Σ ( Xi- X̅ )2
8 9
4 1
10 25
0 25
3 4
Total 25 64

Mean X̅ = Σ Xi / n
X̅ = 25 / 5 = 5


Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 64 / 5 -1 ) = 4

Part b)

Z = ( X - µ ) / σ
Z = ( 8 - 5 ) / 4
Z = 0.75

Z = ( X - µ ) / σ
Z = ( 4 - 5 ) / 4
Z = -0.25

Z = ( X - µ ) / σ
Z = ( 10 - 5 ) / 4
Z = 1.25

Z = ( X - µ ) / σ
Z = ( 0 - 5 ) / 4
Z = -1.25

Z = ( X - µ ) / σ
Z = ( 3 - 5 ) / 4
Z = -0.5